### v1

@@ Abstract
Autoregressive models are often employed to learn distributions of image data by decomposing the $D$-dimensional density function into a product of one-dimensional conditional distributions. Each conditional depends on preceding variables (pixels, in the case of image data), making the order in which variables are processed fundamental to the model performance. In this paper, we study the problem of observing a small subset of image pixels (referred to as a pixel patch) to predict the unobserved parts of the image. As our prediction mechanism, we propose a generalized and computationally efficient version of the convolutional neural autoregressive distribution estimator (ConvNADE) model adapted for real-valued and color images. Moreover, we investigate the quality of image reconstruction when observing both random pixel patches and low-discrepancy pixel patches inspired by quasi-Monte Carlo theory. Experiments on benchmark datasets demonstrate that choosing the pixels akin to a low-discrepancy sequence reduces test loss and produces more realistic reconstructed images.
@@
Autoregressive model, image reconstruction, machine learning, low-discrepancy
@@ 1.1 Autoregressive Models
The transition of autoregressive models from traditional statistics to machine learning occurred as researchers recognized their potential for modeling complex, high-dimensional data. These models leverage the chain rule of probability, which allows them to factorize the joint distribution of all the variables in a distribution into a product of conditional distributions. Specifically, the probability of each variable is modeled as a function of the preceding variables. Frey [7] developed the first autoregressive model of this kind by considering the use of logistic regression models for these conditionals. Later, [1] extended the conditionals to be more flexible by using a single-layer feed-forward neural network with a form of parameter sharing paving the way for more powerful and flexible approaches. This work laid the groundwork for models like the neural autoregressive distribution estimator (NADE) developed in [13], which was introduced as an efficient model primarily due to the advantage of weight sharing across all the conditional probabilities, given an arbitrary ordering of the input variables. NADE’s introduction was significant and influential; it provided a way to model high-dimensional distributions using a neural network that was both computationally efficient and capable of capturing complex dependencies in the data. This approach inspired further developments, such as RNADE for real-valued data [20] and ConvNADE for image data [19], addressing specific limitations of earlier models. The mathematical details of NADE and ConvNADE models are given in Section 2.
@@ 1.2 The Ordering of Input Variables
In [21], the authors recognized the significance of ordering and subsequently introduced the idea of order-agnostic NADE models, i.e., a model not dependent on a specific variable ordering, $o$. One such model, known as DeepNADE, trains the model not on one specific sequence of input variables but by exposing the model to multiple random orderings of the input vector of observations during the training process. This approach encourages the model to generalize across different possible sequences of variable dependencies, leading to more robust performance. Subsequently in [19], it was noted that in the context of image data, the DeepNADE model flattens the two-dimensional image to a one-dimensional vector, hence removing the spatial topology that contains useful information regarding which pixels likely depend on each other. To mitigate this issue and to ensure that the spatial information is not lost during the training process, the ConvNADE model was introduced as a further extension of DeepNADE and employs convolutional layers (as in convolutional neural networks) to preserve the spatial topology of the image. Importantly, while the order-agnostic training that occurs in both ConvNADE and DeepNADE enhances the flexibility and robustness during the inference stage, it also introduces unwanted additional computational complexity during the training stage as the model must effectively learn from a much larger space of possible variable input sequences.
One primary application of autoregressive modeling, and the main subject of this paper, is image reconstruction in the form of inpainting. That is, given $P$ pixels from a $D>P$ pixel image, predict the missing $D-P$ pixels. In this work, our prediction mechanism will be a variant of the ConvNADE model adapted specifically for this image reconstruction task. Namely, instead of conditioning on one pixel at a time in the autoregressive procedure, we will instead condition on a small subset of pixels. This training and reconstruction procedure is explained more fully in Section 3.2.
There are various alternative approaches to address the image reconstruction problem, including image transformers [3, 10] and advanced denoising autoencoders [22, 23], such as the masked autoencoder [9], which are commonly used for pre-training models in transfer learning to enhance image classification tasks [6, 9]. However, these models demand extensive computational resources, often requiring days of training even with state-of-the-art GPU hardware. As a result, they fall outside the scope of this work, which aims to provide a computationally efficient solution for image reconstruction.
@@ 1.3 Our Contribution and Paper Overview
The main goal of this paper is to investigate the quality of full image reconstruction having observed only a subset of the total pixels. Our tool for reconstruction is a generalized convolutional NADE model, ConvNADE-Beta-Color, adapted for the first time for real-valued and color images in Section 3.1. We tailor the training procedure for ConvNADE-Beta-Color specifically for the image reconstruction task in Section 3.2. We show in Section 4 on several popular benchmark datasets including Binarized MNIST, CIFAR-10, FER2013, and LHQ that a observing a low-discrepancy pixel patch yields a more geometric perspective of the entire image when compared to a random patch, yielding a better model fit and more reliable image reconstructions.
@@ 2.3 Convolutional NADE (ConvNADE)
$p(x_{o_{d}}=1|\mathbf{x}_{o_{<d}})=\textbf{vec}(\mathbf{H}^{(L)})_{o_{d}},$
where the notation $\textbf{vec}(\mathbf{X})\rightarrow\mathbf{x}$ refers to concatenation of every row in $\mathbf{X}$ and $\mathbf{H}^{(L)}$ is the output feature of the final convolutional layer. Similarly to DeepNADE, ConvNADE uses a binary mask $\mathbf{M}_{o_{<d}}$, applied element-wise to the input
$\mathbf{H}^{(0)}=\mathbf{X}\odot\mathbf{M}_{o_{<d}},$
@@ 3 A Generalized ConvNADE Model
A limitation of the original NADE and ConvNADE models is their reliance on Bernoulli distributions for conditional probabilities, restricting their applicability to binary datasets. For problems involving real-valued response distributions, the RNADE model [20] is commonly employed, utilizing a mixture of Gaussian distributions to model the response variable. However, this approach requires training a substantial number of parameters – twice the number of mixture components – which makes the model computationally expensive to fit. As highlighted in [20], while Gaussian mixtures are a popular choice for modeling continuous responses, alternative distributions may better suit specific tasks.
To address these limitations, we extend the ConvNADE model discussed in Section 2 to handle color datasets, broadening its utility. Moreover, we demonstrate that using a single Beta distribution to model the response variable is not only feasible but also capable of producing clear, realistic outputs. This approach is significantly more computationally efficient requiring the training of only two parameters, the shape and scale of the Beta distribution.
@@ 3.1 ConvNADE-Beta-Color
where $c=1,2,3$ corresponds to the red, green, and blue color channels and $H^{(L-1)}$ denotes the number of channels in layer $L-1$. The weight matrix in the $L^{th}$ layer is denoted $\mathbf{W}^{(L)}\in\mathbb{R}^{H^{(L-1)}\times 6\times N_{\mathbf{X}}\times N_{\mathbf{X}}}$ is and $b^{(L)}\in\mathbb{R}^{6}$ is bias vector in the $L^{th}$ layer. The final conditional distribution for each channel is then modeled as
@@
for each $c=1,2,3$. Figure 1 presents a simple schematic of the ConvNADE-Beta-Color model where we ensure the first half of its layers uses “valid” convolutions while the other half uses “full” convolutions. It was found that including the mask $\mathbf{M}_{o{<d}}$ as an input to the model improves performance [19], as it resolves ambiguity about whether a value is 00 or has been masked to 00. For ConvNADE-Beta-Color, the mask was provided as an additional channel to the input layer hence the four channel input in Figure 1.
Figure 1: ConvNADE and ConvNADE-Beta-Color with $L=8$ convolutional layers. The number of channels for a given layer is given by the number before the “@” symbol, and after this symbol the filter size and the type of convolution is specified in parentheses. For ConvNADE the image has 2 channels: one for each binary pixel value and one for the mask. The output channel has 1 layer to represent the the Bernoulli parameter. For ConvNADE-Beta-Color the image has 4 channels: one for each color channel and one for the mask. The output channel has 6 layers representing shape and scale parameters for each color.
@@ 3.2 Training ConvNADE-Beta-Color for Image Reconstruction
As previously noted, ConvNADE models are trained on a random subset of pixel orderings sampled from the full set of possible orderings, $\mathcal{D}!$. During training, the model processes one pixel at a time according to a given ordering, where conditioning on each individual pixel is computationally burdensome. On the other hand, several practical applications require the reconstruction of images from a specified or preselected patch of pixels. For example, tasks like image inpainting (otherwise known as image reconstruction) involve predicting and filling in missing pixels based on the given surrounding context [2, 14]. In this section, we propose a method to train the ConvNADE-Beta-Color model to leverage a patch of input pixels for reconstructing the remaining pixels in an image, making it well-suited for this application.
@@
In the following, we describe two pixel patches that will be investigated. A random pixel patch of size $P$ is defined as a subset $\mathcal{S}^{\text{rand}}_{P}\subseteq\{1,\ldots,D\}$ sampled uniformly at random from the set of all subsets of $\{1,\ldots,D\}$, i.e.,
$\mathcal{S}^{\text{rand}}_{P}\sim U\big{(}\{S\subseteq\{1,\ldots,D\}:|S|=P\}\big{)}.$
Geometrically, this corresponds to selecting $P$ pixels uniformly at random without replacement from the entire pixel grid.
@@
Figure 2: A $32\times 32$ image from LHQ illustrating pixel patches. A random patch (Left) and low-discrepancy patch (Right) of size 128.
@@ 4.2 Datasets
We evaluate our models on several classical datasets; Binarized MNIST, FER2013, CIFAR-10, and LHQ. Binarized MNIST serves as a classical benchmark containing images of handwritten digits from 00 to $9$. We use the standard dataset split, consisting of 50000 training examples, 10000 validation examples, and 10000 test examples.
The FER2013 dataset contains centered greyscale facial images with varying expressions. Originally sized at $48\times 48$ pixels, these images are resized to $32\times 32$ pixels to facilitate training. The dataset is divided into 24403 training examples, 4306 validation examples, and 3589 test examples.
The CIFAR-10 dataset includes $32\times 32$ color images of 10 object classes, such as trucks, dogs, and ships, with 6000 images per class. For our experiments, we use a train-validation-test split of 44000 training examples, 6000 validation examples, and 10000 test examples.
Finally, the LHQ dataset consists of colored landscape images originally sized at $256\times 256$ pixels, which are scaled down to $32\times 32$ pixels for training purposes. The dataset is divided into 63000 training examples, 13500 validation examples, and 13500 test examples.
Patches Bin. MNIST FER2013 CIFAR-10 LHQ
Random $0.1104$ $-1.2517$ $-1.0771$ $-1.4562$
Low-discrepancy $\mathbf{0.0994}$ $\mathbf{-1.3367}$ $\mathbf{-1.1536}$ $\mathbf{-1.5468}$
Table 1: The test loss results for the ConvNADE model on the Binarized MNIST test dataset and the ConvNADE-Beta-Color model on FER2013, CIFAR-10, and LHQ datasets.
@@ 4.3 Numerical Results
Table 1 presents the test loss values for both the ConvNADE and ConvNADE-Beta-Color models when trained on random and low-discrepancy pixel patches across several datasets. For the random patches, the valued returned is the mean average loss over five random patches.
@@ 5 Discussion
In this paper, we propose a variant of the convolutional neural autoregressive distribution estimator, referred to as ConvNADE-Beta-Color, designed specifically for reconstructing images from a limited subset of observed pixels. This model extends the original ConvNADE framework to accommodate continuous and color images, significantly broadening its applicability, while having the benefit of being much less computationally expensive than its original counterparts. Through experiments across multiple benchmark datasets, we analyze the impact of different input pixel patches. Our results consistently show that ConvNADE-Beta-Color produces coherent reconstructions and further, observing a low-discrepancy pixel patch leads to lower test loss, faster convergence, and clearer image reconstructions versus observing a random pixel patch. Future work could explore extending this approach to color image generation and applying low-discrepancy pixel patches to enhance image compression tasks.
Appendix A Image reconstructions
This single appendix contains the image reconstructions for all experiments detailed in Section 4.

### v2

@@ Abstract
Autoregressive models are often employed to learn distributions of image data by decomposing the $D$-dimensional density function into a product of one-dimensional conditional distributions. Each conditional depends on preceding variables (pixels, in the case of image data), making the order in which variables are processed fundamental to the model performance. In this paper, we study the problem of observing a small subset of image pixels (referred to as a pixel patch) to predict the unobserved parts of the image. As our prediction mechanism, we propose a generalized version of the convolutional neural autoregressive distribution estimation (ConvNADE) model adapted for real-valued and color images. Moreover, we investigate the quality of image reconstruction when observing both random pixel patches and low-discrepancy pixel patches inspired by quasi-Monte Carlo theory. Experiments on benchmark datasets demonstrate that, where design permits, pixels sampled or stored to preserve uniform coverage improves reconstruction fidelity and test performance.
@@
Autoregressive models, image reconstruction, machine learning, low-discrepancy sampling
@@ 1.1 Autoregressive Models
The transition of autoregressive models from traditional statistics to machine learning occurred as researchers recognized their potential for modeling complex, high-dimensional data. These models leverage the chain rule of probability, which allows them to factorize the joint distribution of all the variables in a distribution into a product of conditional distributions. Specifically, the probability of each variable is modeled as a function of the preceding variables. Frey [10] developed the first autoregressive model of this kind by considering the use of logistic regression models for these conditionals. Later, [2] extended the conditionals to be more flexible by using a single-layer feed-forward neural network with a form of parameter sharing paving the way for more powerful and flexible approaches. This work laid the groundwork for models like the neural autoregressive distribution estimation (NADE) model developed in [17], which was introduced as an efficient model primarily due to the advantage of weight sharing across all the conditional probabilities, given an arbitrary ordering of the input variables. NADE’s introduction was significant and influential; it provided a way to model high-dimensional distributions using a neural network that was both computationally efficient and capable of capturing complex dependencies in the data. This approach inspired further developments, such as RNADE for real-valued data [25] and ConvNADE for image data [24], addressing specific limitations of earlier models. The mathematical details of NADE and ConvNADE models are given in Section 2.
@@ 1.2 The Ordering of Input Variables
In [26], the authors recognized the significance of ordering and subsequently introduced the idea of order-agnostic NADE models, i.e., a model not dependent on a specific variable ordering, $o$. One such model, known as DeepNADE, trains the model not on one specific sequence of input variables but by exposing the model to multiple random orderings of the input vector of observations during the training process. This approach encourages the model to generalize across different possible sequences of variable dependencies, leading to more robust performance. Subsequently in [24], it was noted that in the context of image data, the DeepNADE model flattens the two-dimensional image to a one-dimensional vector, hence removing the spatial topology that contains useful information regarding which pixels likely depend on each other. To mitigate this issue and to ensure that the spatial information is not lost during the training process, the ConvNADE model was introduced as a further extension of DeepNADE that employs convolutional layers (as in convolutional neural networks) to preserve the spatial topology of the image. Importantly, while the order-agnostic training that occurs in both ConvNADE and DeepNADE enhances the flexibility and robustness during the inference stage, it also introduces additional computational complexity during the training stage as the model must effectively learn from a much larger space of possible variable input sequences.
1.3 Image Reconstruction
One primary application of autoregressive modeling, and the main subject of this paper, is image reconstruction. That is, given $P$ pixels from a full $D>P$ pixel image, predict the missing $D-P$ pixels. In this work, our prediction mechanism will be a variant of the ConvNADE model adapted specifically for this image reconstruction task. Namely, instead of conditioning on one pixel at a time in the autoregressive procedure, we will instead condition on a small subset of pixels. This training and reconstruction procedure is explained more fully in Section 3.2.
There are various alternative approaches to address the image reconstruction problem, including image transformers [4, 14] and advanced denoising autoencoders [27, 28], such as the masked autoencoder [13], which are commonly used for pre-training models to enhance image classification tasks [8, 13, 29]. Diffusion models have also become one of the most popular machine learning models for their superiority in image related tasks [6]. However, all of the models mentioned above demand extensive computational resources, and therefore often build upon pre-trained models requiring hundreds of hours of training with hundreds of GPUs distributed across clusters [19, 29]. As a result, they fall outside the scope of this work, which instead aims to provide a more computationally efficient solution for image reconstruction. While our method does not claim to match the reconstruction quality of these high-capacity state-of-the-art models, it offers a lightweight alternative that can feasibly be trained from scratch on a single GPU. Our proposed model also provides a practical framework for investigating how different choices of conditioning pixel subsets influence reconstruction outcomes.
@@ 1.4 Our Contribution and Paper Overview
The main goal of this paper is to investigate the quality of full image reconstruction given only a subset of observed pixels—a setting that arises in several tasks, including image compression when only a small number of pixels are selected for storage [1]. Our tool for reconstruction is a generalized convolutional NADE model, ConvNADE-Beta-Color, which we adapt for the first time to handle real-valued and color images (Section 3.1). We also tailor the training procedure specifically for the image reconstruction task (Section 3.2). In particular, although we build upon the ConvNADE architecture, we depart from the traditional autoregressive formulation by conditioning on a structured subset of observed pixels. Our method is best described as a masked conditional density estimation model built on NADE principles as described above. In Section 4, we evaluate this approach on several benchmark datasets, including Binarized MNIST, CIFAR-10, FER2013, and LHQ. We investigate whether our NADE-inspired model, when conditioned on a low-discrepancy pixel patch (see Section 3.2), can achieve realistic reconstructions while remaining more efficient to train than modern high-capacity alternatives.
@@ 2.3 Convolutional NADE (ConvNADE)
$p(x_{o_{d}}=1|\mathbf{x}_{o_{<d}})=\textbf{vec}(\mathbf{H}^{(L)})_{o_{d}}$
where the notation $\textbf{vec}(\mathbf{X})$ refers to concatenation of every row in $\mathbf{X}$, and $\mathbf{H}^{(L)}$ is the output feature of the final convolutional layer. Similarly to DeepNADE, ConvNADE uses a binary mask $\mathbf{M}_{o_{<d}}$, applied element-wise to the input
$\mathbf{H}^{(0)}=\mathbf{X}\odot\mathbf{M}_{o_{<d}}$
@@ 3 A Generalized ConvNADE Model
A limitation of the original NADE and ConvNADE models is their reliance on Bernoulli distributions for conditional probabilities, restricting their applicability to binary datasets. For problems involving real-valued response distributions, the RNADE model [25] is commonly employed, utilizing a mixture of Gaussian distributions to model the response variable. However, this approach requires training a substantial number of parameters—twice the number of mixture components—which could make the model expensive to fit. As highlighted in [25], while Gaussian mixtures are a popular choice for modeling continuous responses, alternative distributions may better suit specific tasks.
In this text, we extend the ConvNADE model discussed in Section 2 to handle color datasets, broadening its utility. Moreover, we demonstrate that using a single Beta distribution to model the response variable is not only feasible but also capable of producing clear, realistic outputs. This approach is more computationally efficient requiring the training of only two parameters, the shape and scale of the Beta distribution, however, despite not seen in practice with any of the Section 4 experiments, we acknowledge that the uni-modal nature of the Beta distribution could limit its expressiveness when the true conditional distribution is multi-modal.
@@ 3.1 ConvNADE-Beta-Color
Figure 1: ConvNADE and ConvNADE-Beta-Color with $L=8$ convolutional layers, each defined as channels @ filter size (conv type). For ConvNADE, the image has 2 channels: one for the binary pixel value and one for the mask, and outputs a 1 layer Bernoulli parameter map. For ConvNADE-Beta-Color, the image has 4 channels: one for each color and one for the mask and the output channel has $6$ layers representing Beta shape and scale for each color.
@@
where $c=1,2,3$ corresponds to the red, green, and blue color channels and $H^{(L-1)}$ denotes the number of channels in layer $L-1$. The weight matrix in the $L^{th}$ layer is denoted $\mathbf{W}^{(L)}\in\mathbb{R}^{H^{(L-1)}\times 6\times N_{\mathbf{X}}\times N_{\mathbf{X}}}$ and $b^{(L)}\in\mathbb{R}^{6}$ is the bias vector in the $L^{th}$ layer. The final conditional distribution for each channel is then modeled as
@@
for each $c=1,2,3$. Figure 1 presents a simple schematic of the ConvNADE-Beta-Color model where we ensure the first half of its layers uses “valid” convolutions while the other half uses “full” convolutions. It was found that including the mask $\mathbf{M}_{o_{<d}}$ as an input to the model improves performance [24], as it resolves ambiguity about whether a value is 00 or has been masked to 00. For ConvNADE-Beta-Color, the mask was provided as an additional channel to the input layer hence the four channel input in Figure 1.
@@ 3.2 Training ConvNADE-Beta-Color for Image Reconstruction
As previously noted, ConvNADE models are traditionally trained on a random subset of pixel orderings sampled from the full set of permutations, $\mathcal{D}!$. This corresponds to conditioning on one pixel at a time in a fully autoregressive manner. However, in many practical applications, such as image inpainting and data compression, the ‘order’ of pixels is not explicitly known and it may be realistic to condition on a fixed subset of observed pixels [3, 18, 12]. In this section, we depart from the standard NADE objective of modeling the full joint distribution via a chain of one-dimensional conditionals. Instead, we propose a method to train the ConvNADE-Beta-Color model that learns to predict missing pixels given a pre-defined subset. While the architecture, including its use of masking, remains recognizable as derived from its NADE parent, the training procedure is adapted for fixed-pattern conditioning rather than full autoregression.
Figure 2: A $32\times 32$ image from LHQ illustrating pixel patches. (Left) A random patch and (Right) a low-discrepancy patch of size 128.
@@
In the following, we describe two pixel patches that will be investigated. A random pixel patch of size $P$ is defined as a subset $\mathcal{S}^{\text{rand}}_{P}\subseteq\{1,\ldots,D\}$ sampled uniformly at random from the set of all subsets of $\{1,\ldots,D\}$, i.e., $\mathcal{S}^{\text{rand}}_{P}\sim U\big{(}\{S\subseteq\{1,\ldots,D\}:|S|=P\}\big{)}.$ Geometrically, this corresponds to selecting $P$ pixels uniformly at random without replacement from the entire pixel grid.
@@ 4.2 Datasets
We evaluate our models on several classical datasets; Binarized MNIST, FER2013, CIFAR-10, and LHQ. Binarized MNIST serves as a classical benchmark containing images of handwritten digits from 00 to $9$. We use the standard dataset split, consisting of 50000 training examples, 10000 validation examples, and 10000 test examples. The FER2013 dataset contains centered greyscale facial images with varying expressions. Originally sized at $48\times 48$ pixels, these images are resized to $32\times 32$ pixels to facilitate training. The dataset is divided into 24403 training examples, 4306 validation examples, and 3589 test examples. The CIFAR-10 dataset includes $32\times 32$ color images of 10 object classes, such as trucks, dogs, and ships, with 6000 images per class. For our experiments, we use a train-validation-test split of 44000 training examples, 6000 validation examples, and 10000 test examples. Finally, the LHQ dataset consists of colored landscape images originally sized at $256\times 256$ pixels, which are scaled down to $32\times 32$ pixels for training purposes. The dataset is divided into 63000 training examples, 13500 validation examples, and 13500 test examples.
(a) Binarized MNIST
(b) FER2013
(c) CIFAR-10
(d) LHQ
@@ 4.3 Numerical Results
Table 1 presents the test loss values for both the ConvNADE and ConvNADE-Beta-Color models when trained on random and low-discrepancy pixel patches across several datasets. For the random patches, the value returned is the mean average loss over five random patches with $95\%$ confidence intervals provided.
@@
We also compared training times of ConvNADE-Beta-Color with ConvNADE-MoG-Color on the FER2013 dataset. ConvNADE-MoG-Color is otherwise identical to ConvNADE-Beta-Color, as described in Section 3.1, except that it uses a mixture of Gaussians (in the case of this experiment, a mixture of $10$ Gaussians) to model the output distribution for each color channel. We can report that ConvNADE-Beta-Color completed training in 2375 seconds, faster than ConvNADE-MoG-Color, which required 3221 seconds across 120 epochs. In addition to this improved efficiency, ConvNADE-Beta-Color avoids the need to select the number of mixture components, a hyperparameter required by ConvNADE-MoG-Color.
Patches Bin. MNIST FER2013
Random $0.11$ $(0.102,0.119)$ $-1.252$ $(-1.257,-1.247)$
LD $\mathbf{0.0994}$ $\mathbf{-1.3367}$
Patches CIFAR-10 LHQ
Random $-1.077$ $(-1.086,-1.068)$ $-1.456$ $(-1.484,-1.429)$
LD $\mathbf{-1.1536}$ $\mathbf{-1.5468}$
Table 1: Test loss results for the ConvNADE model on the Binarized MNIST dataset, and for the ConvNADE-Beta-Color model on FER2013, CIFAR-10, and LHQ datasets. Results for random patches are mean average loss over 5 random patches with $95\%$ CI.
@@ 5 Discussion
In this paper, we propose a variant of the convolutional neural autoregressive distribution estimator, referred to as ConvNADE-Beta-Color, designed specifically for reconstructing images from a limited subset of observed pixels. This model extends the original ConvNADE framework to accommodate continuous and color images, significantly broadening its applicability, while having the benefit of being much less computationally expensive than its original counterparts. Through experiments across multiple benchmark datasets, we analyze the impact of different input pixel patches. Our results consistently show that ConvNADE-Beta-Color produces coherent reconstructions and further, observing a low-discrepancy pixel patch leads to lower test loss, faster convergence, and clearer image reconstructions versus observing a randomly selected patch.
Although in many practical settings the observed pixels are fixed by the acquisition process, our results suggest that whenever there is flexibility in which pixels are observed, choosing them with uniformity is beneficial. This principle is relevant in systems where pixel access can be controlled such as single-pixel cameras [9] and superpixel-based image compression [1] where the observation pattern can be explicitly designed.
Future work could explore extending this approach to color image generation and applying low-discrepancy pixel patches to enhance image compression tasks.
Appendix A Image reconstructions from Section 4 Experiments